# Provably Convergent Plug-and-play Proximal Block Coordinate Descent Method for Hyperspectral Anomaly Detection

### v1

@@ 4.3 Subsequence convergence
Combining the inequalities above, we obtain (26) with $c_{1}=\min\{\delta-\rho_{1}+\alpha_{\mathcal{S}},\alpha_{E},\delta-\rho_{2}+\alpha_{\mathcal{Z}}\}$.
(ii) Since $(E^{k})^{\top}E^{k}={I}_{r}$, we have the sequence $\{E^{k}\}$ is bounded. By (i), we have $F(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\leq F(\mathcal{Z}^{0},E^{0},\mathcal{S}^{0})$. Also, we observe that $F(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\geq\tau\|\mathcal{S}^{k}\|_{2,\psi}+\Phi_{\Sigma}(\mathcal{Z}^{k})\geq 0$. Since both $\|\cdot\|_{2,\psi}$ and $\Phi_{\Sigma}$ are coercive, that is,
@@
we have the sequences $\{\mathcal{S}^{k}\}$ and $\{\mathcal{Z}^{k}\}$ are bounded.
@@
Assume that assumptions (A1) and (A2) are satisfied. Let $\{(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\}$ be the sequence generated by Algorithm 1 with $\alpha_{\mathcal{Z}}>\rho_{1}-\delta$ and $\alpha_{\mathcal{S}}>\rho_{2}-\delta$. Then $\lim_{k\to\infty}\|\mathcal{Z}^{k}\times_{3}E^{k}-\mathcal{Z}^{k-1}\times_{3}E^{k-1}\|_{F}=0$.
@@
Assume that assumptions (A1) and (A2) are satisfied. Let $\{(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\}$ be the sequence generated by Algorithm 1 with $\alpha_{\mathcal{Z}}>\rho_{1}-\delta$ and $\alpha_{\mathcal{S}}>\rho_{2}-\delta$. Then every accumulation point of $\{(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\}$ is a first-order stationary point of problem (2).
@@ 5 Numerical Experiments
To evaluate the accuracy of object detection, we plot the receiver operating characteristic (ROC) curve [15] and calculate the area under the ROC curve (AUC) [16]. The ROC curve plots the probability of detection vs the false alarm rate for various possible thresholds. The closer the ROC curve is to the upper-left corner and the larger AUC score, the better the detection performance. The average AUC scores and average computational time in seconds are presented in Table 1. The ROC curves obtained from HSIs with no noise and with a noise level of $0.03$ are presented in Figure 2 and Figure 3, respectively. As shown in Table 1, when no noise is presented, the proposed PnP-PBCD method achieves the highest average AUC score. The other methods have lower average AUC scores due to poor performance in one or more scenes, as shown in Figure 2. This issue becomes more visible when the observed HSI is degraded by noise. For example, the LRASR method performs competitively with our method in “Urban 1” and “Urban 2” as shown in Figure 3 (c) and (d). However, it performs poorly for “Airport” and “Beach” as shown in Figure 3 (a) and (b). Hence, the proposed PnP-PBCD method is more robust in anomaly detection across different test HSIs, even with noise.
@@
All in all, the proposed PnP-PBCD method is an efficient method for anomaly detection, outperforming the competing methods both numerically and visually, especially in noisy HSIs.

### v3

@@ 4.2 Subsequence convergence
Combining the inequalities above, we obtain (22) with $c_{1}=\min\{\alpha_{\mathcal{S}}+(\alpha_{\mathcal{S}}+\delta-\rho_{1})_{+},\alpha_{E},\alpha_{\mathcal{Z}}+(\alpha_{\mathcal{Z}}+\delta-\rho_{2})_{+}\}$.
(ii) Since $(E^{k})^{\top}E^{k}={I}_{r}$, we have that the sequence $\{E^{k}\}$ is bounded. By (i), we have $F(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\leq F(\mathcal{Z}^{0},E^{0},\mathcal{S}^{0})$. Also, we observe that $F(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\geq\tau\|\mathcal{S}^{k}\|_{2,\psi}+\Phi_{\Sigma}(\mathcal{Z}^{k})\geq 0$. Since both $\|\cdot\|_{2,\psi}$ and $\Phi_{\Sigma}$ are coercive, that is,
@@
we have that the sequences $\{\mathcal{S}^{k}\}$ and $\{\mathcal{Z}^{k}\}$ are bounded.
@@
Assume that assumptions (A1) and (A2) are satisfied. Let $\{(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\}$ be the sequence generated by Algorithm 1 with $\alpha_{\mathcal{S}}+(\alpha_{\mathcal{S}}+\delta-\rho_{1})_{+}>0$ and $\alpha_{\mathcal{Z}}+(\alpha_{\mathcal{Z}}+\delta-\rho_{2})_{+}>0$. Then $\lim_{k\to\infty}\|\mathcal{Z}^{k}\times_{3}E^{k}-\mathcal{Z}^{k-1}\times_{3}E^{k-1}\|_{F}=0$.
@@
Assume that assumptions (A1) and (A2) are satisfied. Let $\{(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\}$ be the sequence generated by Algorithm 1 with $\alpha_{\mathcal{S}}+(\alpha_{\mathcal{S}}+\delta-\rho_{1})_{+}>0$ and $\alpha_{\mathcal{Z}}+(\alpha_{\mathcal{Z}}+\delta-\rho_{2})_{+}>0$. Then every accumulation point of $\{(\mathcal{Z}^{k},E^{k},\mathcal{S}^{k})\}$ is a first-order stationary point of problem (1).
@@ 5 Numerical Experiments
5.1 Comparison of experimental results
To evaluate the accuracy of object detection, we plot the receiver operating characteristic (ROC) curve [15] and calculate the area under the ROC curve (AUC) [16]. The ROC curve plots the probability of detection vs the false alarm rate for various possible thresholds. The closer the ROC curve is to the upper-left corner and the larger AUC score, the better the detection performance.
The average AUC scores and average computational time in seconds are presented in Table 1. The ROC curves obtained from HSIs with no noise and with a noise level of $0.03$ are presented in Figure 2 and Figure 3, respectively. In terms of computational time, as shown in Table 1, the proposed PnP-PBCD method completes detection in an average of less than 15 seconds, although the traditional RX method remains the fastest overall. Regarding detection accuracy, as also shown in Table 1, when no noise is presented, the proposed PnP-PBCD method achieves the highest average AUC score. The other methods have lower average AUC scores due to poor performance in one or more scenes, as shown in Figure 2. This issue becomes more visible when the observed HSI is degraded by noise. For example, the LRASR method performs competitively with our method in “Urban 1” and “Urban 2” as shown in Figure 3 (c) and (d). However, it performs poorly for “Airport” and “Beach” as shown in Figure 3 (a) and (b). Hence, the proposed PnP-PBCD method is more robust in anomaly detection across different test HSIs, even with noise.
@@ 5.1 Comparison of experimental results
All in all, the proposed PnP-PBCD method is an efficient method for anomaly detection, outperforming the competing methods in terms of detection accuracy and visual quality, especially in noisy HSIs.
@@
5.2 Parameter settings and analysis
For the proposed PnP-PBCD method, we set model parameters $\delta=0.25$, and $\tau=1$; and we select the relaxed $\ell_{p}$ norm with $p=0.1$ and $\varepsilon=10^{-5}$ as $\psi$ for $\|\cdot\|_{2,\psi}$ given in (2). We also set the parameters $a=0.2$, $b=0.4$, and $\gamma=0.99$ for the shifted and relaxed denoiser $\widetilde{\mathcal{D}}_{\sigma}^{\gamma}$ given in (5); and we set algorithm parameters $\alpha_{\mathcal{S}}=\alpha_{E}=\alpha_{\mathcal{Z}}=0.01$. Algorithm 1 stops if $\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|_{F}/\|\mathcal{S}^{k}\|_{F}\leq 10^{-3}$.
Next, we conduct a parameter analysis on the model parameters $\delta$ and $\tau$. We use Urban 1 with a noise level of $0.03$ for testing. The plot of AUC values vs parameters $\delta$ and $\tau$ is presented in Fig. 8. The AUC values are similar when the ratio of $\delta$ to $\tau$ (i.e., $\frac{\delta}{\tau}$) is fixed, and the AUC achieves the best performance when this ratio is approximately $0.25$.
Figure 8: Plot of the AUC values vs parameters $\delta$ and $\gamma$.